6. And finally someone who just heard acid, decides to use H2SO4 (since it’s the only one they can remember. So, they put 5.0g in 1.0 liters of solution. What pH do you now expet? (Remember that sulfuric acid only loses 1 H+, so it’s concentration is the value of H3O+.)
1
Expert's answer
2020-06-08T15:08:55-0400

Normality of 1L solution with 5g of H2SO4:

N= m(H2SO4)*2/Mr(H2SO4)= 5*2/98=0.1 N

pH=-lg(0.1)=1;

Answer: The ph of the solution is 1.